OpenStudy (anonymous):

$5400  is invested, part of it at  11%  and part of it at  6%.  For a certain year, the total yield is  $ 474.00.   How much was invested at each rate? @Directrix

OpenStudy (johnweldon1993):

So think of it as a system of equations We know total there was 5400 invested into 2 different accounts 'x' and 'y' so \[\large x + y = 5400\] Now we know that if 'x' has an 11% interest rate and 'y' has a 6% interest rate...and together those interests totaled 474..we have \[\large .11x + .06y = 474\] Now just solve that system

OpenStudy (anonymous):

so I would 5400* .11 and 5400 * 0.6?

OpenStudy (johnweldon1993):

Not quite, do you know how to solve a system of equations?

OpenStudy (anonymous):

yea...a little...That's what I am learning in school as we speak

OpenStudy (anonymous):

which method would I use? sub or elimination?

OpenStudy (johnweldon1993):

Well I would just use substitution here.. In that first equation, solve for either 'x' or 'y' ... your choice

OpenStudy (anonymous):

I suck at substitution

OpenStudy (johnweldon1993):

Alright, so \[\large x + y = 5400\] Do you see how I got this equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so I plug in x and y as 5400

OpenStudy (johnweldon1993):

And to check, do you see where I got the second equation as well?

OpenStudy (anonymous):

yea

OpenStudy (johnweldon1993):

Alright good, back to the first equation \[\large x + y = 5400\] we cannot say that x and y are 5400....that wouldnt make sense...instead, we can solve or 1 of those variables...lets say 'x' To solve this equation for 'x'...lets subtract 'y' from both sides of the equation \[\large x = 5400 - y\] make sense?

OpenStudy (anonymous):

yea

OpenStudy (johnweldon1993):

With that, and the fact that we have a second equation \[\large .11x + .06y = 474\] Since, from the first equation,we now know what 'x' equals...we can "substitute" it into this equation for 'x' So if \[\large x = 5400 - y\] then this second equation will now be \[\large .11(5400 - y) + 0.6y = 474\] see what happened there?

OpenStudy (anonymous):

yea..u plugged it in

OpenStudy (johnweldon1993):

Right, Now we just solve...that .11 outside the parenthesis will be distributed...multiplied into both terms inside the parenthesis...so \[\large .11(5400 - y) + 0.6y = 474\] will now become \[\large 594 - .11y + 0.6y = 474\] We can simplify that further, and solve for 'y'...can you do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

make y by itself?

OpenStudy (johnweldon1993):

Well, yes we can start with that...we have 2 'y' terms...how could we get them by themselves on 1 side of the equation?

OpenStudy (anonymous):

subtract 11y from both sides

OpenStudy (johnweldon1993):

Nope, alright so \[\large 594 - .11y + 0.6y = 474\] is what we have Lets start by combining the 2 'y' terms... what is -0.11y + 0.6y?

OpenStudy (anonymous):

11.6

OpenStudy (johnweldon1993):

No, we have 2 decimals \[\large -0.11y + 0.6y=?\]

OpenStudy (anonymous):

0.71...my bad

OpenStudy (johnweldon1993):

Remember that the 0.11 is NEGATIVE so we are adding the 0.6 to a negative 0.11

OpenStudy (anonymous):

-0.49

OpenStudy (johnweldon1993):

So close, but still not quite...

OpenStudy (johnweldon1993):

Try it the other way what is 0.6 minus 0.11 ?

OpenStudy (anonymous):

0.49